The length of a rectangular plot of land exceeds its breadth by 23m.if the length is decreased by 15m and the breadth is increased by 7m, the area is reduced by 360m.m.find the length and breadth of the plot?

Answer by Boreal(15235)

Length plot 1=x
Width plot 1=x-23
A=x(x-23)=x^2-23x
;
Length plot 2: =x-15
Width plot 2:x-23+7= x-16
Area plot 2=(x-15)(x-16)=x^2-31x+240
But plot 2 is 360 m^2 less than plot 1
They equal each other if 360 is added to plot 2
Therefore, x^2-23x=x^2-31x+600
x^2 cancel
8x=600
x=75 m
x-23=52 m
Plot 1 is 75 x 52=3900 m^2
Plot 2 is 60 + 59=3540 m^2, which is 360 less than 3900.
The answer is plot 1, 75 m long and 52 m wide.